OpenStudy (anonymous):

a(t)=(2,0,6t) V(0)=(0,1,0) r(0)=(0,1,-3) how to find unit tangent and unit normal vector when t=1

OpenStudy (anonymous):

tangent and normal require knowledge of V9t) and\[v(t) = v(0) + a*t\]

OpenStudy (anonymous):

i got T=(2t,1,3t^2)/(4t^2+1+9t^4)^(1/2) then i am stuck

OpenStudy (anonymous):

I was wrong. here\[v(t)= v(0) + \int\limits_0^t a(t) dt\]

OpenStudy (anonymous):

But you made a wrong computation

OpenStudy (anonymous):

\[v(t) = (0,1,0) + (2t, 0, 3t^2)\]

OpenStudy (anonymous):

v(t) =(2t,1,3t^2) am i right??

OpenStudy (anonymous):

substitute t = 1

OpenStudy (anonymous):

unit tangent = v(t)/Iv(t)I

OpenStudy (anonymous):

so i got (2t,1,3t^2)/(4t^2+1+9t^4)^(1/2)

OpenStudy (anonymous):

It is clearly unneeded in general form. 1 Substitute 2 Compute || and divide by it

OpenStudy (anonymous):

Do I need a equation to compute unit normal vector?

OpenStudy (anonymous):

Just 3 numbers that you obtain by substitution of t=1 into the expression v(t)

OpenStudy (anonymous):

i think i need a general form of unit tangent to compute unit normal vector N=T'/IT'I

OpenStudy (anonymous):

No just the |T| at one point.

OpenStudy (anonymous):

thx i think u are right

OpenStudy (anonymous):

close q.